OpenStudy (anonymous):

use implicit differentiation to find an equation of the tangent of the curve at the given point. ( will post the problem in a second)

OpenStudy (anonymous):

\[2(x ^{2}+y ^{2})^{2}=25(X ^{2}+y ^{2}) \] the point is (3,1)

OpenStudy (zale101):

\[\frac{ d }{ dx }[2(x^{2} + y^{2})^2]=25 \frac{ d }{ dx }[(x^2+y^2]\] The chain rule should be used in the left side of the equation \[4(x^2+y^2)(x^2+y^2)'\]=\frac{ d }{ dx }[25(x^2+y^2]\] for the left side of the equation, we could use the power rule for x^2 and the chanin rule for y^2 4(2x+2y(y)'=25(2x+2y(y)')

OpenStudy (zale101):

\[4(x^2+y^2)(2x+2y(y)')=25(2x+2y(y)')\]

OpenStudy (zale101):

Now, solve for y prime (y)' :)

OpenStudy (zale101):

Did i make it clear? :)

OpenStudy (anonymous):

i think so. I am trying to solve for y prime real fast

OpenStudy (zale101):

Ok ^.^

OpenStudy (anonymous):

I have no idea. I keep trying but its not coming out right. can you divide both sides by 2x+2y(yprime)?

OpenStudy (anonymous):

which would leave you with no yprimes? allowing you to solve for y and get your tangent line?

OpenStudy (zale101):

I'll just work the problem through to see if you got it right \[4(x^2+y^2)(2x+2y(z))=25(2x+2yz)\] Think of y prime as z, and you are trying to solve for it \[=(8x^2+4y^2)(2x+2yz)=50x+50yz\] expand (8x^2+4y^2)(2x+2yz) \[=8x^3+8xy^2+8x^2yz+8y^3z=50x+50yz\] subtract 50yz from both sides \[=-50y+8x^2y^2-50yz+8x^2yz+8y^3z=50x\] factor z out \[=(-50y+8x^2y+8y^3)z=50x-8x^3-8xy^2\] now, divide everything, except z to the left side of the equation to isolate z by itself \[z=\frac{ (50x-8x^3-8xy^2) }{ -50y+8x^2y+8y^3 }\]

OpenStudy (anonymous):

i wasnt factoring.. that might have been my first issue haha

OpenStudy (zale101):

Yeah, haha factoring is really important when differentiating algebra is the main tool in calculus, if something was wrong, blame it on the algebra part :)

OpenStudy (anonymous):

alright I think i got it! and thanks!

OpenStudy (zale101):

No problem, and good luck!